# Radial Transport in High-Redshift Disk Galaxies Dominated by Inflowing Streams

### v1

@@ 5.3 Stream and Non-Stream Signal: Statistics
Most panels show a large scatter in the plotted quantities. Nonetheless, the median trends are illuminating. For the metallicity-based selection, the mass in streams (non-streams) increases (decreases) with increasing $r$, with the median mass attaining a value close to $\mathscr{M}$ (zero) near $r=R_{\rm d}$. In the streamline-based selection, the scatter is larger. However, the median stream (non-stream) mass increases (decreases) with increasing $r$, attains a maximum (minimum) of $\sim\!0.4\ (0.6)\ \mathscr{M}$ around $\sim\!0.4\ R_{d}$, and then slightly decreases (increases) to $\sim\!0.3\ (0.7)\ \mathscr{M}$ near $r=R_{\rm d}$. Note that in the metallicity-based selection, the mass in streams (non-streams) dominates at large (small) $r$, with the median stream mass being greater than $0.5\mathscr{M}$ for $r\gtrsim 0.5\ R_{\rm d}$. In the streamline-based selection, however, the mass in non-streams is greater than in streams (at least at the median level) at all $r$. Nevertheless, in both selections, the median non-stream flux is close to zero, and the median cold gas flux at large $r$ ($r\gtrsim 0.5\ R_{\rm d}$ in the metallicity-based selection and $r\gtrsim 0.3\ R_{\rm d}$ in the streamline-based selection) is dominated by the streams. At smaller $r$, the median stream flux tends to zero and is smaller than the median non-stream flux, such that the median cold gas flux, although very small, is dominated by the non-streams. The median stream $V_{r}$ is almost always negative, with typically larger magnitudes in the streamline-based selection than in the metallicity-based selection (the same is also true for the stream flux). The median non-stream $V_{r}$ is typically close to zero and smaller in magnitude than the median stream $V_{r}$, especially at large $r$. The median cold gas $V_{r}$ is always negative (also discussed in §3.2) but generally smaller in magnitude than the median stream $V_{r}$. This is expected as the cold gas $V_{r}$ is a mass-weighted average of the stream and non-stream $V_{r}$. From the median trends in $F_{r}$ and $V_{r}$, we conclude that, on average, the radial transport of cold gas in the VELA disk galaxy sample at large $r$ is dominated by the inflowing motion of recently accreted streams.
@@
From the middle panels, we note that the inflowing and outflowing non-stream fluxes in the streamline-based selection are not negligible (see also §5.2). In fact, they are similar in magnitude to the inflowing stream flux in the same selection. However, due to their comparable magnitude, they almost cancel each other, yielding a total non-stream flux that is close to zero at the median level, albeit with a large scatter (top-middle panel of Fig. 11). The outflowing stream flux in this selection is negligible, as the selection procedure does not generally identify outflowing regions in the galaxy (which are also likely to be outflowing for the past dynamical time) as streams (see §5.2). The inflowing and outflowing non-stream fluxes in the metallicity-based selection are also of comparable magnitude and can cancel each other. However, even by themselves, they are negligible compared to the stream flux in the same selection, except at small $r$ (inside $\sim\!0.5\ R_{\rm d}$), where the stream flux is also small. This is because of the increasingly smaller mass in non-streams than in streams with increasing $r$. As such, the non-stream flux in this selection is also close to zero, with very little scatter (top-middle panel of Fig. 11). The outflowing stream flux in the metallicity-based selection is not negligible, as in this selection, outflowing regions may get identified as streams based on their metallicity. However, we note that the outflowing stream flux is typically smaller in magnitude than the inflowing stream flux such that the total stream flux (top-left panel of Fig. 11) is mostly negative.
@@
Figure 13: Comparison of Observations with the VELA Disk Galaxy Sample: $\langle V_{r}\rangle/\langle V_{\rm rot}\rangle$ is plotted against $f$ for all cold gas (top panel), streams (middle and bottom left-hand panels), and non-streams (middle and bottom right-hand panels), with each circle corresponding to a particular snapshot in the VELA disk galaxy sample, color-coded by its redshift, $z$. Here, $\langle V_{r}\rangle$ is the stream-, non-stream-, or disk-averaged radial velocity, as relevant, $\langle V_{\rm rot}\rangle$ is the disk-averaged rotational velocity, and $f$ is the cold gas fraction. The red squares and the vertical error bars denote the medians and $16^{\rm th}-84^{\rm th}$ percentiles, respectively, in $f$ bins of varying widths, as indicated by the horizontal error bars, with the locations of the squares along the $f$-axis being at the bin-averages. The results from the metallicity- and streamline-based selections are shown in the middle and bottom rows, respectively. The green diamonds with vertical error bars highlight the Genzel et al. (2023) observations. The observations agree with the stream-averaged radial velocities from the simulations (middle and bottom left-hand panels). The disk-averaged (top row) and non-stream-averaged (middle and bottom right-hand panels) radial velocities from the simulations typically have much smaller magnitudes and barely agree with the observations.
@@ 6 Comparison with Observations
Therefore, instead of using the disk-instability-driven transport models, here, we interpret the Genzel et al. (2023) observations in light of the VELA simulations. In Figs. 13 and 14, $\langle V_{r}\rangle/\langle V_{\rm rot}\rangle$ is plotted against the cold gas fraction, $f$, and redshift, $z$, respectively, for all cold gas (top panel), streams (middle and bottom left-hand panels), and non-streams (middle and bottom right-hand panels), with each circle corresponding to a particular snapshot in the VELA disk galaxy sample, color-coded by $z$ in Fig. 13 and by $f$ in Fig. 14. Recall that $\langle V_{r}\rangle$ is the stream-, non-stream-, or disk-averaged radial velocity, as relevant, and $\langle V_{\rm rot}\rangle$ is the disk-averaged rotational velocity. The red squares and the vertical error bars denote the medians and $16^{\rm th}-84^{\rm th}$ percentiles, respectively, in $f$ and $z$ bins of varying widths, as indicated by the horizontal error bars, with the locations of the squares along the $f$- and $z$-axes being at the bin-averages. The results from the metallicity- and streamline-based selections are shown in the middle and bottom rows, respectively. The green diamonds with vertical error bars highlight the Genzel et al. (2023) observations.
The Genzel et al. (2023) galaxies are located at redshifts between $1$ to $2.5$ and have large gas fractions ($f>0.35$). On the other hand, our disk galaxy sample spans a redshift range of $\sim\!1$ to $\sim\!6$ and has cold gas fractions varying from $\sim\!0.01$ to $\sim\!0.7$ (most of the sample has $f<0.35$). Nevertheless, where they overlap in $z$ and $f$, the Genzel et al. (2023) observations agree with the stream-averaged radial velocities from the simulations (middle and bottom left-hand panels). Note that the stream-averaged radial velocities are largely similar in both selections, capable of attaining large negative values (up to $\sim\!0.6-0.8$ times the disk-averaged rotational velocity). One exception is that the stream-averaged radial velocity in the streamline-based selection is always negative, whereas that in the metallicity-based selection can occasionally be positive (up to $\sim\!0.2$ times the disk-averaged rotational velocity). The disk-averaged (top rows) and non-stream-averaged (middle and bottom right-hand panels) radial velocities from the simulations typically have much smaller magnitudes and barely agree with the observations.
@@
We thus conclude that the inflow velocities within the streams in the simulated disks may explain the radial velocities deduced from the observations if the observed signal is dominated by recently accreted streams and is characteristic of their (mass-weighted) average radial motions. A more robust comparison in the “observational plane” is yet to be performed, where mock observations of the simulated disks are analyzed the same way as the observed galaxies. Apart from accretion along cold streams, gas transport along a bar may be another way to achieve large radial motions (Pastras et al. in preparation). However, only one of the nine galaxies in the Genzel et al. (2023) sample shows strong evidence for the presence of a bar, and it is unclear whether bar-driven transport is also important for the rest of the sample.
@@ 7 Conclusion
    Using two different crude stream selection techniques, one based on metallicities and another on streamlines, the cold gas within the disk is classified into recently accreted streams versus off-stream material. On average, the streams are found to dominate the radial inflow outside the inner disk ($r\gtrsim 0.3R_{\rm d}$ for the streamline-based selection and r $\gtrsim 0.5R_{\rm d}$ for the metallicity-based selection).
@@
We conclude that the inward radial transport of cold gas in extended, star-forming disk galaxies at cosmic noon is dominated by the external radial velocities of the incoming streams from the cosmic web rather than being driven by torques from internal disk instabilities. Future studies with smoothed-particle hydrodynamics simulations or AMR simulations with tracer particles, where recently accreted gas along cold streams can be identified with better accuracy, should further solidify our novel result.
Our finding that the disk-averaged inward radial transport in the cosmological simulations is significantly weaker than the predictions of the analytical models for isolated disks poses an open question for future study. One option is that the dominant presence of anisotropic incoming streams from the cosmic web changes the overall conditions in the disks in a way that invalidates some of the assumptions made in the analytical models. In particular, it may affect the energy balance in the turbulent disk between streams, feedback, dissipation, and radial transport. Also, the premise that $Q$ is self-regulated to a value less than or equal to unity may not be valid for the simulated disks. Indeed, by analyzing a few of the VELA disks, Inoue et al. (2016) found that while collapsed clumps tend to have $Q<1$, proto-clump regions generally have $Q\gtrsim 2-3$. The high $Q$-values possibly represent excessive compressive modes of turbulence, likely induced by external perturbations, such as intense accretion along the incoming streams (see also Mandelker et al., 2024, Ginzburg et al. in preparation). However, there are also indications from various simulations of isolated disks for low averaged radial inflows, with significant deviations from axial symmetry, qualitatively similar to the findings from our cosmological simulations (private communications with Andreas Burkert and Oscar Agertz). This indicates that the weaknesses in the analytical models based on disk instability are more profound than missing the incoming streams.

### v2

@@ 5.4 Stream and Non-Stream Signal: Statistics
Most panels show a large scatter in the plotted quantities. Nonetheless, the median trends are illuminating. For the metallicity-based selection, the mass in streams (non-streams) increases (decreases) with increasing $r$, with the median mass attaining a value close to $\mathscr{M}$ (zero) near $r=R_{\rm d}$. In the streamline-based selection, the scatter is larger. However, the median stream (non-stream) mass increases (decreases) with increasing $r$, attains a maximum (minimum) of $\sim\!0.4\ (0.6)\ \mathscr{M}$ around $\sim\!0.4\ R_{d}$, and then slightly decreases (increases) to $\sim\!0.3\ (0.7)\ \mathscr{M}$ near $r=R_{\rm d}$. Note that in the metallicity-based selection, the mass in streams (non-streams) dominates at large (small) $r$, with the median stream mass being greater than $0.5\mathscr{M}$ for $r\gtrsim 0.5\ R_{\rm d}$. In the streamline-based selection, however, the mass in non-streams is greater than in streams (at least at the median level) at all $r$. Nevertheless, in both selections, the median non-stream flux is close to zero, and the median cold gas flux at large $r$ ($r\gtrsim 0.5\ R_{\rm d}$ in the metallicity-based selection and $r\gtrsim 0.3\ R_{\rm d}$ in the streamline-based selection) is dominated by the streams. At smaller $r$, the median stream flux tends to zero and is smaller than the median non-stream flux, such that the median cold gas flux, although very small, is dominated by the non-streams. The median stream $V_{r}$ is almost always negative, with typically larger magnitudes in the streamline-based selection than in the metallicity-based selection (the same is also true for the stream flux). The median non-stream $V_{r}$ is typically close to zero and smaller in magnitude than the median stream $V_{r}$, especially at large $r$. The median cold gas $V_{r}$ is always negative (also discussed in §3.2) but generally smaller in magnitude than the median stream $V_{r}$. This is expected as the cold gas $V_{r}$ is a mass-weighted average of the stream and non-stream $V_{r}$. From the median trends in $F_{r}$ and $V_{r}$, we conclude that, on average, the radial transport of cold gas in the VELA disk galaxy sample at large $r$ is dominated by the inflowing motion of recently accreted streams. These trends are also robust against the details of the feedback physics employed in the simulations, which we have verified by comparing these results with a newer generation of the VELA simulations (Ceverino et al., 2023) that uses the same set of initial conditions but has stronger stellar feedback (see Appendix A).
@@
From the middle panels, we note that the inflowing and outflowing non-stream fluxes in the streamline-based selection are not negligible (see also §5.3). In fact, they are similar in magnitude to the inflowing stream flux in the same selection. However, due to their comparable magnitude, they almost cancel each other, yielding a total non-stream flux that is close to zero at the median level, albeit with a large scatter (top-middle panel of Fig. 12). The outflowing stream flux in this selection is negligible, as the selection procedure does not generally identify outflowing regions in the galaxy (which are also likely to be outflowing for the past dynamical time) as streams (see §5.3). The inflowing and outflowing non-stream fluxes in the metallicity-based selection are also of comparable magnitude and can cancel each other. However, even by themselves, they are negligible compared to the stream flux in the same selection, except at small $r$ (inside $\sim\!0.5\ R_{\rm d}$), where the stream flux is also small. This is because of the increasingly smaller mass in non-streams than in streams with increasing $r$. As such, the non-stream flux in this selection is also close to zero, with very little scatter (top-middle panel of Fig. 12). The outflowing stream flux in the metallicity-based selection is not negligible, as in this selection, outflowing regions may get identified as streams based on their metallicity. However, we note that the outflowing stream flux is typically smaller in magnitude than the inflowing stream flux such that the total stream flux (top-left panel of Fig. 12) is mostly negative. We also note that the non-stream inflowing and outflowing fluxes could just be non-circular motions that average out to zero. As we define inflows and outflows by the sign of the radial velocity, non-circular motions would be considered as flows in our analysis, and we can’t distinguish one from the other.
The marked difference in the stream mass in the two different selections, as seen in the bottom panel of Figure 12, warrants some discussion. It is possible that the streamline-based selection is more reliable in the inner disk, due to increasing metallicities towards the galaxy centers, which may make it difficult to distinguish between streams and non-streams in the metallicity-based selection (see also Section 5.3 and Figs. 9 and 10). But even the streamline-based selection shows declining stream mass inside $\sim 0.2\ R_{{\rm d}}$, albeit not as stark as the decline in the metallicity-based selection. Over a few dynamical times, the streams are likely to phase-mix with the rest of the disk, and may become indistinguishable from the rest of the disk. However, more robust stream identification using tracer particles is required to understand the survival of streams within disks both temporally and spatially.
@@
Figure 14: Comparison of Observations with the VELA Disk Galaxy Sample: $\langle V_{r}\rangle/\langle V_{\rm rot}\rangle$ is plotted as a function of the cold gas fraction ($f$) for all cold gas (top panel), streams (middle and bottom left-hand panels), and non-streams (middle and bottom right-hand panels). Each circle represents a snapshot from the VELA disk galaxy sample and is color-coded by its corresponding redshift, $z$. Here, $\langle V_{r}\rangle$ denotes the stream-, non-stream-, or disk-averaged radial velocity, as relevant, and $\langle V_{\rm rot}\rangle$ is the disk-averaged rotational velocity. Red squares indicate the median values within bins of $f$, with vertical error bars showing the $16^{\rm th}$$84^{\rm th}$ percentile ranges and horizontal bars reflecting the bin widths. The horizontal position of each square corresponds to the average $f$ within that bin. Results from the metallicity-based and streamline-based classifications are shown in the middle and bottom rows, respectively. For comparison, observational data from Genzel et al. (2023) are represented by green diamonds with vertical error bars. These observations are consistent with the stream-averaged radial velocities obtained from the simulations (middle and bottom left panels). In contrast, the disk-averaged (top row) and non-stream-averaged (middle and bottom right panels) radial velocities are generally of much smaller magnitude and show limited agreement with the observed values.
@@
Figure 16: The left- and right-hand panels show the histograms for cold gas fraction, $f$, and redshift, $z$, for the VELA disks in red and the Genzel et al. (2023) galaxies in green. The VELA disks span a wider redshift range and have typically lower gas fractions than the Genzel et al. (2023) sample. While the lower gas fractions in VELA could potentially bias our conclusions, the simulated sample does include a few gas rich galaxies. But the disk-averaged radial velocities in even these gas-rich galaxies barely reproduce the observations, which are much better explained when assumed to be originating from recently accreted streams.
@@ 6 Comparison with Observations
Therefore, instead of using the disk-instability-driven transport models, here, we interpret the Genzel et al. (2023) observations in light of the VELA simulations. In Figs. 14 and 15, $\langle V_{r}\rangle/\langle V_{\rm rot}\rangle$ is plotted against the cold gas fraction, $f$, and redshift, $z$, respectively, for all cold gas (top panel), streams (middle and bottom left-hand panels), and non-streams (middle and bottom right-hand panels), with each circle corresponding to a particular snapshot in the VELA disk galaxy sample, color-coded by $z$ in Fig. 14 and by $f$ in Fig. 15. Recall that $\langle V_{r}\rangle$ is the stream-, non-stream-, or disk-averaged radial velocity, as relevant, and $\langle V_{\rm rot}\rangle$ is the disk-averaged rotational velocity. The red squares and the vertical error bars denote the medians and $16^{\rm th}-84^{\rm th}$ percentiles, respectively, in $f$ and $z$ bins of varying widths, as indicated by the horizontal error bars, with the locations of the squares along the $f$- and $z$-axes being at the bin-averages. The results from the metallicity- and streamline-based selections are shown in the middle and bottom rows, respectively. The green diamonds with vertical error bars highlight the Genzel et al. (2023) observations, where $f$ corresponds to molecular gas fraction.
In Fig 16, the left- and right-hand panels show the histograms for $f$ and $z$ for the Genzel et al. (2023) galaxies in green and the VELA disk galaxy sample in red. The Genzel et al. (2023) galaxies are located at redshifts between $1$ to $2.5$ and have large gas fractions ($f>0.35$). On the other hand, our disk galaxy sample spans a redshift range of $\sim\!1$ to $\sim\!6$ and has cold gas fractions varying from $\sim\!0.01$ to $\sim\!0.7$ (most of the sample has $f<0.35$). Nevertheless, where they overlap in $z$ and $f$, the Genzel et al. (2023) observations agree with the stream-averaged radial velocities from the simulations (middle and bottom left-hand panels of Figures 14 and 15). Note that the stream-averaged radial velocities are largely similar in both selections, capable of attaining large negative values (up to $\sim\!0.6-0.8$ times the disk-averaged rotational velocity). One exception is that the stream-averaged radial velocity in the streamline-based selection is always negative, whereas that in the metallicity-based selection can occasionally be positive (up to $\sim\!0.2$ times the disk-averaged rotational velocity). The disk-averaged (top rows) and non-stream-averaged (middle and bottom right-hand panels) radial velocities from the simulations typically have much smaller magnitudes and barely agree with the observations.
@@
[…]
@@ 7 Conclusion
    Using two different crude stream selection techniques, one based on metallicities and another on streamlines, the cold gas within the disk is classified into recently accreted streams versus off-stream material. On average, the streams are found to dominate the radial mass flux outside the inner disk ($r\gtrsim 0.3R_{\rm d}$ for the streamline-based selection and r $\gtrsim 0.5R_{\rm d}$ for the metallicity-based selection).
@@
We conclude that the inward radial transport of cold gas in extended, star-forming disk galaxies at cosmic noon is dominated by the external radial velocities of the incoming streams that penetrate the halo from the cosmic web and accrete onto the disk rather than being driven by torques from internal disk instabilities. Future studies with smoothed-particle hydrodynamics simulations or AMR simulations with tracer particles, where recently accreted gas along cold streams can be identified with better accuracy, should further solidify our novel result. We note that while the current ART code does not have the capability to add tracers, AMR simulations with tracers can be performed using other codes such as RAMSES (Teyssier, 2002).
Our finding that the disk-averaged inward radial transport in the cosmological simulations is significantly weaker than the predictions of the analytical models for isolated disks poses an open question for future study. One option is that the dominant presence of anisotropic incoming streams from the cosmic web changes the overall conditions in the disks in a way that invalidates some of the assumptions made in the analytical models. In particular, it may affect the energy balance in the turbulent disk between streams, feedback, dissipation, and radial transport. Also, the premise that $Q$ is self-regulated to a value less than or equal to unity may not be valid for the simulated disks. Indeed, by analyzing a few of the VELA disks, Inoue et al. (2016) found that while collapsed clumps tend to have $Q<1$, proto-clump regions generally have $Q\gtrsim 2-3$. The high $Q$-values possibly represent excessive compressive modes of turbulence, likely induced by external perturbations, such as intense accretion along the incoming streams (see also Mandelker et al., 2024; Ginzburg et al., 2025). However, there are also indications from simulations of isolated disks for low averaged radial inflows, with significant deviations from axial symmetry, qualitatively similar to the findings from our cosmological simulations (private communications with Andreas Burkert and Oscar Agertz). This may indicate weaknesses in the analytical models based on disk instability more profound than missing the incoming streams. It is also possible that contribution to gas turbulence-driving in such isolated disks by stellar feedback reduces the radial inflow rates compared to that predicted by the various disk-instability-based models ignoring feedback (Krumholz et al., 2018). However, previous work by Goldbaum et al. (2016) using isolated disk simulations with gas fractions upto $20\%$ show that including feedback only modestly lowers the levels of radial inflow compared to that in simulations without feedback (Goldbaum et al., 2015).